Manisha is 5 years younger than meghana.four years later,meghana will be twice as old as manisha.find their present ages.
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52754)

.
Let x be Manisha's present age.

Then Meghana's present age is (x+5) years.

Four years later: Manisha is (x+4). Meghana is (x+5) + 4 = (x+9).

The condition says 

2*(x + 4) = x + 9.

2x + 8 = x + 9.

x = 1.


Answer.  Manisha is 1 year old now. Meghana present age is 1 + 5 = 6 years.

Solved.

----------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".



Answer by greenestamps(13195)


Let's solve this using logical reasoning instead of formal algebra.

For the sake of word problems involving ages, we can assume that Manisha will ALWAYS be 5 years younger than Meghana. So 4 years from now Meghana will still be 5 years older than Manisha; if she is then twice as old as Manisha, their ages 4 years from now must be 10 and 5.

So their current ages are 6 and 1.